In an experiment similar to the one you are about to perform, a student
found that the molarity of acetic acid in a vinegar was 0.640 mol/L.
(a) Calculate the acetic acid in g/L.
(b) What is the concentration of acetic acid in g per 100 mL of vinegar?
(c) What, therefore, is the weight by volume percentage (w/v %) of
acetic acid in the vinegar?
(d) Is the product information on the la
1
Expert's answer
2019-09-06T04:15:03-0400

In an experiment similar to the one you are about to perform, a student

found that the molarity of acetic acid in a vinegar was 0.640 mol/L

Solution:

(a) Calculate the acetic acid in g/L.

C(vinegar) = M(vinegar)*c(vinegar)

C(vinegar) = 60g/mol * 0.640 mol/L = 38.4 g/L

(b) What is the concentration of acetic acid in g per 100 mL of vinegar?

C(vinegar) = 3.84 g per 100 mL.

(c) What, therefore, the weight by volume percentage (w/v %) of a acetic acid in the vinegar?

% w/v = g of solute/100 mL of solution

%w/v = 3.84%

(d) Is the product information on the la

The product information have the concentration of acetic acid in vinegar as a v/v%:

v/v % =volume of acetic acid

volume of vinegar× 100.